OpenStudy (anonymous):

2.15 g of an unknown gas at 71 degrees C and 1.10 atm is stored in a 1.75 L flask. What is the density of the gas? What is the molar mass of the gas? 2.15 g of an unknown gas at 71 degrees C and 1.10 atm is stored in a 1.75 L flask. What is the density of the gas? What is the molar mass of the gas? @Chemistry

OpenStudy (matt101):

Density is m/V = 2.15/1750 = 0.00123 g/mL Use the universal gas law, PV=nRT, to find the molar mass (MM). Note that n = m/MM, so PV=(m/MM)RT --> MM = (mRT)/(PV) MM = [(2.15)(0.08206)(344)]/[(1.1)(1.75)] = 31.53 g/mol Remember to convert temperature to Kelvin.

OpenStudy (anonymous):

Thanks!